9. Use Maxwells speed distribution law to derive the expression for the most probable speed of molecules in a sample of gas RTdv
Answer #1

Given That

\frac{dN}{N} = \frac{4}{\sqrt{\pi }}\left ( \frac{M}{2RT} \right )^{3/2}v^{2}e^{\frac{-Mv^{2}}{2RT}}dv

dN = \frac{4N}{\sqrt{\pi }}\left ( \frac{M}{2RT} \right )^{3/2}v^{2}e^{\frac{-Mv^{2}}{2RT}}dv

To get most probable speed we have to find maxima i.e putting this derivative equal to zero

i.e.

\frac{4N}{\sqrt{\pi }}\left ( \frac{M}{2RT} \right )^{3/2}v^{2}e^{\frac{-Mv^{2}}{2RT}}dv = 0

Here constant can not be zero i.e

\frac{d\left (v^{2}e^{\frac{-Mv^{2}}{2RT}} \right )}{dv} = 0

Using chain rule

0 = 2v.e^{\frac{-Mv^{2}}{2RT}}+v^{2}.\left (\frac{-2Mv}{2RT} \right )e^{\frac{-Mv^{2}}{2RT}}

Cancelling common terms

0 = 1-v^{2}.\left (\frac{M}{2RT} \right )

v^{2}.\left (\frac{M}{2RT} \right ) = 1


v^{2} = \frac{2RT}{M}

Here M = N.m

v^{2} = \frac{2RT}{N.m} R/N = kB


v^{2} = \frac{2k_{B}T}{m}


v_{mp} =\sqrt{ \frac{2k_{B}T}{m}}